# Existence of All Wilton Ripples of the Kawahara Equation

### v1

@@ Abstract
The existence of all small-amplitude Wilton ripple solutions of the Kawahara equation is proven. These are periodic, traveling-wave solutions that bifurcate from a two-dimensional nullspace spanned by two distinct, co-propagating cosine waves. In contrast with previous results, the proof, which relies on a carefully constructed Lyapunov-Schmidt reduction, implies the existence of all small-amplitude Wilton ripples of the Kawahara equation, of which there are countably infinite. Though this result pertains only to the Kawahara equation, the method of proof likely extends to most nonlinear dispersive equations admitting Wilton ripple solutions.
@@
Acknowledgments: This work was funded by the National Science Foundation through grant NSF-DMS-2402044.
@@ 1 Introduction
We consider the Kawahara (or super-KdV) equation
@@
where $\alpha,\beta,$ and $\sigma$ are real, nonzero parameters. First derived in [13], the Kawahara equation models shallow water gravity-capillary waves near a critical value of the Bond number, a nondimensional quantity that compares gravity and surface tension. Owing to the competition between these forces, the Kawahara equation admits several interesting and physically-relevant solutions; see, for instance, [4]-[5] and [15, 21, 25]. Of importance in this work is a class of periodic, traveling-wave solutions called the Wilton ripples, which have a particularly rich history in the context of water waves [^1] More recently, these solutions have found practical relevance in the context of magnetohydrodynamics [26]. [24].
To motivate the mathematical behavior of Wilton ripples, we begin by searching for the periodic, traveling-wave solutions of the Kawahara equation. To this effect, we move to a traveling frame $x\rightarrow x-ct$ so that (1) becomes
@@
for all $\kappa>0$. In fact, by rescaling variables according to $x\rightarrow\kappa^{-1}x$, $u\rightarrow\alpha k^{2}\sigma^{-1}u$, $c\rightarrow\alpha k^{2}c$, and $\beta\rightarrow\alpha\kappa^{-2}\beta$, we can instead investigate the $2\pi$-periodic solutions of
$cu+u_{xx}+\beta u_{4x}+u^{2}=0,$ (5)
without loss of generality. Furthermore, because (5) is autonomous in $x$ and satisfies the discrete symmetry $x\rightarrow-x$, we need only consider its $2\pi$-periodic, even solutions.
If we further restrict these solutions to small-amplitude, (5) effectively linearizes to
@@
    where $u_{r}(x;a)$ is a smooth, $2\pi$-periodic function in $x$ that is real analytic in $a$ with $||u_{r}(x;a)||_{\textrm{L}^{2}}=\mathcal{O}\left(a^{2}\right)$ as $a\rightarrow 0$ and $c_{r}(a)$ is an even, real-analytic function in $a$ such that $|c_{r}(a)|=\mathcal{O}\left(a^{2}\right)$ as $a\rightarrow 0$. We call such solutions the Stokes waves of the Kawahara equation by analogy with the corresponding solutions of the surface gravity wave equations [22]. Their existence for sufficiently small $a$ was proven in [9] and their stability discussed in [7, 9, 17, 23]. The profile of a typical Stokes wave solution is provided in Figure 1(a).
@@
    When $\beta=1/(1+K^{2})$, the periodic traveling-wave solutions bifurcate from a superposition of two co-propagating cosine waves. These are the Wilton ripples of the Kawahara equation, named by analogy with the corresponding solutions of the gravity-capillary wave equations [24]. Unlike the Stokes waves, Wilton ripples arise from a two-dimensional bifurcation problem, and as a consequence, their expansion for small $a$ is different than that of the Stokes waves. Indeed, based on the formal asymptotic expansions derived in [10, 14], the Wilton ripple solutions of the Kawahara equation take the form
@@
Partial existence proofs have emerged for Wilton ripples of the Kawahara equation and other nonlinear dispersive equations [2]-[3],[8], including those for gravity-capillary water waves [12, 16],[19]-[20]. Nearly all of these proofs, however, restrict to the simplest possible ripple that bifurcates from a superposition of $\cos(x)$ and $\cos(2x)$. In the context of the Kawahara equation, this is the special case when $K=2$.
A notable exception is [8], which adopts a new framework for studying Wilton ripples of a gravity-capillary Whitham equation. In this framework, both $a$ and $b$ are fixed, and the remaining parameters in the governing equation of the Wilton ripples [^3] This would be $\beta$ in the Kawahara equation. become functions of $a$ and $b$. As a consequence, one can prove the existence of a two-dimensional “sheet” of Wilton ripples parameterized by $a$ and $b$. Recent numerical and formal asymptotic investigations by [1] have applied this new framework to the Kawahara equation, and the results suggest that the Wilton ripples discussed herein are embedded in this sheet of solutions (at least for $K=2$ and $K=3$). However, no rigorous proof of this embedding has been established. Thus, the existence of Wilton ripples for all $K\in\mathbb{N}\setminus\{1\}$ remains an open problem. It is the objective of this work to prove this fact. To the author’s knowledge, this is the first proof of its kind to demonstrate the existence of all Wilton ripple solutions of a nonlinear dispersive equation.
In contrast to many previous existence proofs, which rely on Banach fixed point arguments or the method of majorants, this proof relies on a carefully constructed Lyapunov-Schmidt reduction to prove that each Wilton ripple depends analytically on a small-amplitude parameter. It is anticipated that this part of the proof extends with minimal changes to Wilton ripples of any nonlinear dispersive equation, including the gravity-capillary water wave equations. However, to finish the existence proof, one must show that the function $b(a)$ in (8) does not vanish identically for each $K\in\mathbb{N}\setminus\{1\}$. In order to show this, we need high-order asymptotic expansions of all Wilton ripples. For general nonlinear dispersive equations, these expansions have not been carried out, but recent work by [14] has derived these expansions for the Kawahara equation. Using these expansions, we can prove our main result:
@@ 2.1 Notation and Set-Up
$\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T}):=\left\{f:f(-\pi)=f(\pi),~{}f(-x)=f(x),~{}\textrm{and}~{}\sum_{k=0}^{\infty}(1+k^{2})^{4}\left|\widehat{f}_{k}\right|^{2}<+\infty\right\},$ (18)
@@
[…]
@@
[…]
As discussed in the Introduction, $c_{0}+\mathcal{L}$ has a two-dimensional nullspace spanned by $\cos(x)$ and $\cos(Kx)$ when $c_{0}:=1-\beta$. Owing to self-adjointness, the cokernel of $c_{0}+\mathcal{L}$ is also spanned by $\cos(x)$ and $\cos(Kx)$, which guarantees that $c_{0}+\mathcal{L}$ is a Fredholm operator of index zero.
@@
for some $a,b\in\mathbb{R}$, while $u_{r}$ is treated as a small correction to $u_{0}$ involving Fourier cosines modes that are complementary to $\cos(x)$ and $\cos(Kx)$.
@@
We are now ready to proceed with Lyapunov-Schmidt reduction on (26).
@@ 2.2 Solving the Auxiliary Equation
We introduce the projection operators $\mathcal{P}$ and $\mathcal{Q}$ according to
@@
Moreover, since the operator $F$ in (20) is real analytic with respect to $u$ and $c$ and the projections $\mathcal{P}$ and $\mathcal{Q}$ are bounded linear operators, $\mathcal{A}$ is real analytic with respect to $u_{r}$, $a$, $b$, and $c_{r}$. A straightforward calculation shows that
@@
As a consequence, the implicit function theorem gives the existence and uniqueness of
@@
such that, for sufficiently small $a$, $b$, and $c_{r}$, we have (i) $u_{r}(x;a,b,c_{r})\in\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\subset\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$, (ii) $\mathcal{A}(u_{r}(x;a,b,c_{r}),a,b,c_{r})=0$, (iii) $u_{r}(x;0,0,0)=0$, and (iv) $u_{r}(x;a,b,c_{r})$ is jointly real analytic with respect to $a$, $b$, and $c_{r}$.
By analyticity, (34) admits a Taylor series expansion of the form
@@
$\Big{|}\Big{|}R_{3}(x;a,b,c_{r})\Big{|}\Big{|}_{\textrm{H}^{4}(\mathbb{T})}=\mathcal{O}\left((|a|+|b|+|c_{r}|)^{3}\right)\quad\textrm{as}\quad a,b,c_{r}\rightarrow 0;$ (36)
see [6] for details.
@@ 2.3 Solving the Bifurcation Equation
~$\tilde{\mathcal{B}}(a,b,c_{r})=\begin{pmatrix}\mathcal{F}_{1}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]+ac_{r}\\ \mathcal{F}_{K}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]+bc_{r}\end{pmatrix}.$ (40)
@@
Owing to the analyticity of $u_{r}(x;a,b,c_{r})$, we expand $\tilde{\mathcal{B}}$ for sufficiently small $a$, $b$, and $c_{r}$. To this end, we let $\omega\in\{1,K\}$ and consider
@@
$\left|\mathcal{T}_{\omega,3}\right|\leq||u_{r}(x;a,b,c_{r})||^{2}_{\textrm{L}^{2}(\mathbb{T})}\leq||u_{r}(x;a,b,c_{r})||^{2}_{\textrm{H}^{4}(\mathbb{T})}.$ (43)
@@
$||u_{r}(x;a,b,c_{r})||_{\textrm{H}^{4}(\mathbb{T})}\leq C\left(|a|+|b|+|c_{r}|\right)^{2}$ (44)
for some constant $C=C(K)>0$ by Taylor’s theorem. Thus,
@@
regardless of $\omega$. In contrast, $\mathcal{T}_{\omega,1}$ and $\mathcal{T}_{\omega,2}$ contribute terms lower than fourth degree in $a$, $b$, and $c_{r}$ when expanded. The coefficients of these terms are different depending on whether $K=2$, $K=3$, or $K\geq 4$. Thus, we must expand the bifurcation equation in each of these cases separately.

### v2

@@ Abstract
We investigate the existence of Wilton ripple solutions of the Kawahara equation. Without loss of generality, these are $2\pi$-periodic, traveling-wave solutions whose profiles at zero amplitude have a codimension-1 bifurcation from a linear combination of $\cos(x)$ and $\cos(Kx)$ for $K\in\mathbb{N}\setminus\{1\}$. Using a Lyapunov-Schmidt reduction, we prove the existence of these solutions for all $K$, in contrast to previous work demonstrating existence only for $K=2$. Although the proof holds only for the Kawahara equation, many ideas introduced in the proof can be applied to more general contexts, including Wilton ripples of the gravity-capillary water wave equations.
@@
Funding Acknowledgments: This work was funded by the National Science Foundation through grant NSF-DMS-2402044.
Data Availability Statement: The Mathematica notebook CompanionToKawaharaWiltonRipples.nb can be downloaded at https://github.com/rpac5130/KawaharaWiltonRipple.
Conflict of Interest Statement: On behalf of all authors, the corresponding author states that there is no conflict of interest.
@@ 1 Introduction
We consider the Kawahara equation [^1] Alternatively, the fifth-order KdV or super-KdV equation.
@@
where $\alpha,\beta,$ and $\sigma$ are real, nonzero parameters. First derived in [15], the Kawahara equation models shallow water gravity-capillary waves near a critical value of the Bond number, a non-dimensional quantity that compares the forces of gravity and surface tension. In this context, the parameters $\alpha$ and $\sigma$ represent shallow-water dispersion and steepening, as in the KdV equation, while the parameter $\beta$ is a proxy for the strength of surface tension. Because of the inclusion of the surface tension term, the Kawahara equation admits several interesting and physically relevant solutions; see, for instance, [4]-[5] and [18, 24, 31]. Of importance in this work is a class of periodic, traveling-wave solutions called the Wilton ripples.
In this work, we define Wilton ripples as a $2\pi/\kappa$-periodic, traveling-wave solution [^2] Momentarily, we will take $\kappa=1$ without loss of generality. whose profile at zero amplitude has a codimension-1 bifurcation from a linear combination of $\cos(\kappa x)$ and $\cos(K\kappa x)$ for some $K\in\mathbb{N}\setminus\{1\}$. In other words, Wilton ripples represent a special class of periodic traveling waves that are parameterized by one small-amplitude parameter $a$ and have a resonant, two-mode behavior when $a=0$, where the ratio of the modes is $1:K$. Other definitions appear in the literature: some restrict Wilton ripples to the 1:2 resonance only [14, 22, 23, 26], while others view them more broadly as a sheet of solutions for which the codimension of the bifurcation is greater than 1 [9, 16, 19, 29]. Our choice of definition is consistent with that originally intended by Wilton in the context of gravity-capillary water waves [1, 28, 30] and has been used in other fields that study these solutions, including magnetohydrodynamics [32].
To motivate the mathematical behavior of Wilton ripples in the context of the Kawahara equation, we begin by searching for periodic, traveling-wave solutions. To this effect, we move to a traveling frame $x\rightarrow x-ct$ so that (1) becomes
@@
for all $\kappa>0$. In fact, we can rescale $x$ and $u$ according to $x\rightarrow\kappa^{-1}x$ and $u\rightarrow\alpha k^{2}\sigma^{-1}u$, respectively. Then, upon redefining the parameters $c$ and $\beta$ in terms of $\overline{c}=c/(\alpha k^{2})$, and $\overline{\beta}=\kappa^{2}\beta/\alpha$, respectively, we can instead investigate the $2\pi$-periodic solutions of
$\overline{c}u+u_{xx}+\overline{\beta}u_{4x}+u^{2}=0,$ (5)
without loss of generality. In this rescaled equation, only the parameters $\overline{c}$ and $\overline{\beta}$ remain. The parameter $\overline{c}$ retains its physical interpretation as a non-dimensional velocity of the traveling wave, while $\overline{\beta}$ serves as a proxy for the Bond number. More precisely, $\overline{\beta}$ compares the relative magnitude of $\beta/\alpha$ for a fixed periodic wave. Hence, increasing $\overline{\beta}$ increases the overall effect of dispersion due to surface tension over dispersion due to gravitational effects alone. For ease of notation, the overbars on $c$ and $\beta$ will be henceforth removed.
Because (5) is autonomous in $x$ and satisfies the discrete symmetry $x\rightarrow-x$, we need only consider its $2\pi$-periodic, even solutions. If we further restrict these solutions to small-amplitude, (5) effectively linearizes to
@@
    where $u_{r}(x;a)$ is a smooth, $2\pi$-periodic function in $x$ that is real analytic in $a$ with $u_{r}(x;a)||_{\textrm{L}^{2}}=\mathcal{O}\left(a^{2}\right)$ as $a\rightarrow 0$ and $c_{r}(a)$ is an even, real-analytic function in $a$ such that $|c_{r}(a)|=\mathcal{O}\left(a^{2}\right)$ as $a\rightarrow 0$. We call such solutions the Stokes waves of the Kawahara equation by analogy with the corresponding solutions of the surface gravity wave equations [25]. Their existence for sufficiently small $a$ was proven in [11] and their stability discussed in [8, 11, 20, 27]. The profile of a typical Stokes wave solution is provided in Figure 1(a).
@@
    When $\beta=1/(1+K^{2})$, the periodic traveling-wave solutions bifurcate from a superposition of two co-propagating cosine waves with wavenumbers in ratio $1:K$. These are the Wilton ripples of the Kawahara equation, according to our definition. The Wilton ripple solutions have fundamentally different expansions for small $a$ than those of the Stokes waves. Indeed, from the formal asymptotic expansions derived in [12, 17], Wilton ripple solutions of the Kawahara equation take the form
@@
Existence proofs for Wilton ripples have emerged in a variety of settings, beginning with the classical gravity–capillary water wave equations. Reeder and Shinbrot [22, 23] gave the first rigorous construction of small-amplitude Wilton ripples in finite and infinite depth for the 1:21\!:\!2 resonance case using a Lyapunov–Schmidt reduction. Toland and Jones [26, 14] extended these considerations to small-amplitude traveling waves with codimension-1 and -2 bifurcations from flat water involving $\cos(Kx)$ and $\cos((K+1)x)$, of which the 1:21\!:\!2 Wilton ripple is a special case. More recent studies have constructed multimodal resonant solutions (i.e., those with codimension greater than one) in various settings, including gravity–capillary waves with vorticity [9, 16, 19] and stratification [29].
Despite the considerable progress in understanding multimodal traveling waves, the existence of 1:K1\!:\!K Wilton ripple solutions of the gravity–capillary water wave equations remains an open problem for general $K\in\mathbb{N}\setminus\{1\}$. The main challenge lies in tracking certain higher-order terms in the small-amplitude expansion of the Wilton ripples. Specifically, for each $K$, one must determine the order in $a$ at which the $\cos(Kx)$ mode first appears in the expansion. As $K$ increases, this order increases as well, adding to the challenge. For the full gravity–capillary equations, identifying the leading-order contribution of the $\cos(Kx)$ mode for arbitrary $K$ remains intractable. However, for simpler model equations such as the Kawahara equation, the situation is more favorable.
For the Kawahara equation and related weakly nonlinear models, the existence of 1:21\!:\!2 Wilton ripples has been rigorously established via the method of majorants [2, 3]. In principle, this technique can be extended to 1:K1\!:\!K ripples for general $K$, but as noted in [1], the complexity of the required computations grows rapidly with $K$, for essentially the same reasons as in the full water wave setting. An alternate method developed in [10] for the gravity–capillary Whitham equation circumvents this issue by allowing the amplitudes of $\cos(x)$ and $\cos(Kx)$ to serve as free parameters, treating the remaining coefficients in the governing equation [^4] For the gravity–capillary Whitham equation, this would be the Bond number. For the Kawahara equation, this would be $\beta$. as dependent variables. This leads to a construction of two-dimensional sheets of solutions for each $K$, rather than discrete branches. However, since Wilton ripples are traditionally defined for fixed values of equation parameters, it is unclear whether the classical 1:K1\!:\!K branches lie embedded in these sheets. Numerical and formal asymptotic evidence from [1] supports this embedding when $K=2$ and $K=3$ for Wilton ripples of the Kawahara equation and other weakly nonlinear models, but a rigorous justification is still lacking.
Thus, the existence of 1:K1\!:\!K Wilton ripples for all $K\in\mathbb{N}\setminus\{1\}$ remains unresolved, even for model equations. The objective of this work is to close this gap for the Kawahara equation. To the author’s knowledge, this is the first rigorous proof of the existence of all 1:K1\!:\!K Wilton ripple solutions in a nonlinear dispersive PDE. The proof uses Lyapunov–Schmidt reduction to demonstrate analytic dependence of the Wilton ripple solutions on a small-amplitude parameter $a$. The remaining technical obstacle then is to prove that the coefficient $b(a)$ in (8) does not vanish identically for any $K\geq 2$. This requires expansions of the Wilton ripple solutions to arbitrarily high order in $a$, which have only recently been computed for the Kawahara equation in [17] using formal methods. By rigorously justifying these expansions, we prove that $b(a)$, in fact, does not vanish identically and, hence, obtain the existence of the following $1:K$ Wilton ripple solutions of the Kawahara equation:
@@ 2.1 Notation and Set-Up
$\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T}):=\left\{f(x)=\sum_{k=0}^{\infty}\hat{f}_{k}\cos(kx)\quad\textrm{s.t.}\quad\sum_{k=0}^{\infty}(1+k^{2})^{4}\left|\widehat{f}_{k}\right|^{2}<+\infty\right\},$ (18)
@@
$\widehat{f}_{k}=\mathcal{F}_{k}[f(x)]=\begin{cases}\frac{1}{\pi}\int_{-\pi}^{\pi}f(x)\cos(kx)dx&k>0\\ \frac{1}{2\pi}\int_{-\pi}^{\pi}f(x)dx&k=0\end{cases},$ (19)
@@
$\left<v,w\right>=\int_{-\pi}^{\pi}v(x)w(x)dx.$ (22)
As shown in the Introduction, $c_{0}+\mathcal{L}$ has a two-dimensional nullspace spanned by $\cos(x)$ and $\cos(Kx)$ when $c_{0}:=1-\beta$. Owing to self-adjointness, the cokernel of $c_{0}+\mathcal{L}$ is also spanned by $\cos(x)$ and $\cos(Kx)$, which guarantees that $c_{0}+\mathcal{L}$ is a Fredholm operator of index zero.
@@
for some $a,b\in\mathbb{R}$, while $u_{r}$ is treated as a small correction to $u_{0}$ involving Fourier cosines modes orthogonal to $\cos(x)$ and $\cos(Kx)$.
@@
Equation (26) will be the starting point of our Lyapunov-Schmidt reduction.
An intermediate step of the reduction will show the existence of a function $u_{r}\in\textrm{Null}\{c_{0}+\mathcal{L}_{0}\}^{\perp}$ that has real analytic dependence on parameters $a$, $b$, and $c_{r}$. We will require a Taylor series expansion of this $u_{r}$ around $(a,b,c_{r})=(0,0,0)$, and norms for the remainders of this series will be expressed in big-oh notation. For clarity, given any $f\in X$, where $X$ is a Banach space [^5] Typically, $\textrm{H}^{4}_{\textrm{per,even}}(\mathbb{T})$ or $\mathbb{R}^{3}$., and given any $n\in\mathbb{R}^{+}$, we say
[…]
provided there exist $C=C(n)>0$ and $\delta=\delta(n)>0$ such that
[…]
With the necessary notation established, we proceed with Lyapunov-Schmidt reduction on (26).
@@ 2.2 Solving the Auxiliary Equation
We introduce the operators $\mathcal{P}$ and $\mathcal{Q}$ according to
@@
Moreover, since $F$ in (20) is real analytic [^6] Indeed, $\mathcal{L}$ is a bounded linear operator from $\textrm{H}_{\textrm{per,even}}^{4}(\mathbb{T})$ to $\textrm{L}_{\textrm{per,even}}^{2}(\mathbb{T})$, and $u^{2}$ is an analytic operator from $\textrm{H}_{\textrm{per,even}}^{4}(\mathbb{T})$ to $\textrm{L}_{\textrm{per,even}}^{2}(\mathbb{T})$ because $\textrm{H}_{\textrm{per,even}}^{4}(\mathbb{T})$ is a Banach algebra. with respect to $u$ and $c$ and since the projections $\mathcal{P}$ and $\mathcal{Q}$ are bounded linear operators, $\mathcal{A}$ is real analytic with respect to $u_{r}$, $a$, $b$, and $c_{r}$.
Using analyticity, a straightforward calculation shows that
@@
Thus, we have $\mathcal{A}:\textrm{Null}\left(c_{0}+\mathcal{L}\right)^{\perp}\times\mathbb{R}^{3}\rightarrow\textrm{Null}\left(c_{0}+\mathcal{L}\right)^{\perp}$ such that
  1. (i)
    $\textrm{Null}\left(c_{0}+\mathcal{L}\right)^{\perp}$ with the $\textrm{H}^{4}$ norm and $\mathbb{R}^{3}$ with the $\ell^{1}$ norm (for convenience) are Banach spaces,
  2. (ii)
    $\mathcal{A}$ is real analytic with respect to both $u_{r}$ and $(a,b,c_{r})^{T}$,
  3. (iii)
    $\mathcal{A}$ satisfies $\mathcal{A}(0,0,0,0)=0$, and
  4. (iv)
    $\partial_{u_{r}}\mathcal{A}(0,0,0,0)$ is boundedly invertible.
Hence, we can apply the analytic version of the Implicit Function Theorem in Banach spaces [6, Theorem 2.3] to deduce the existence and uniqueness of
@@
such that, for sufficiently small $a$, $b$, and $c_{r}$, we have
  1. (i)
    $u_{r}(x;a,b,c_{r})\in\textrm{Null}\{c_{0}+\mathcal{L}\}^{\perp}\subset\textrm{H}^{4}_{\textrm{per},\textrm{even}}(\mathbb{T})$,
  2. (ii)
    $\mathcal{A}(u_{r}(x;a,b,c_{r}),a,b,c_{r})=0$,
  3. (iii)
    $u_{r}(x;0,0,0)=0$, and
  4. (iv)
    $u_{r}(x;a,b,c_{r})$ is real analytic with respect to $a$, $b$, and $c_{r}$.
As (36) is a real analytic map from $\mathbb{R}^{3}$ to $\textrm{Null}\left(c_{0}+\mathcal{L}\right)^{\perp}$, we can use Taylor’s Theorem in the Banach space setting [7, Theorem 5.4] to expand $u_{r}$ as a multi-variable power series in $a$, $b$, and $c_{r}$ with values in $\textrm{Null}\left(c_{0}+\mathcal{L}\right)^{\perp}$. In particular, we have
@@
[…]
@@ 2.3 Solving the Bifurcation Equation
$\tilde{\mathcal{B}}(a,b,c_{r})=\begin{pmatrix}\mathcal{F}_{1}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]+ac_{r}\\ \mathcal{F}_{K}\left[\left(a\cos(x)+b\cos(Kx)+u_{r}(x;a,b,c_{r})\right)^{2}\right]+bc_{r}\end{pmatrix}$. (42)
@@
Owing to the analyticity of $u_{r}(x;a,b,c_{r})$ with respect to $(a,b,c_{r})^{T}$, we expand $\tilde{\mathcal{B}}$ for sufficiently small $a$, $b$, and $c_{r}$. To this end, we let $\omega\in\{1,K\}$ and consider
@@
[…]
@@
[…]
for some constant $C>0$ by Taylor’s Theorem [^7] Note that this constant must be independent of $a$, $b$, and $c_{r}$, but it may depend on $K$.. Thus,
@@
regardless of $\omega$. In contrast, $\mathcal{T}_{\omega,1}$ and $\mathcal{T}_{\omega,2}$ contribute terms lower than fourth degree in $a$, $b$, and $c_{r}$ when expanded. Ultimately, the differences in the leading-order behavior of $\mathcal{T}_{\omega,1}$ and $\mathcal{T}_{\omega,2}$ determine the number of distinct branches of Wilton ripple solutions when $K=2$, $K=3$, and $K\geq 4$. Indeed, we will find two branches of real-valued solutions to the bifurcation equation when $K=2$ and, hence, two branches of Wilton ripple solutions in this case, while three branches are found when $K=3$ and only one branch is found when $K\geq 4$.